# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ 2.1. Quasi-Cyclic Codes and Their Representations
Let $s\in\mathbb{Z}_{m}$ and let $C_{s}=\{s,sq,sq^{2},...,sq^{r-1}\}\subset\mathbb{Z}_{m}$ be the $q$-cyclotomic coset of $s$, where $r$ is the smallest positive integer so that $sq^{r}\equiv s\mod m$. It is important to note that the $q$-cyclotomic cosets split $\mathbb{Z}_{m}.$ This classic concept is seen in constructions of cyclic codes, as well as in the next theorem (formally known as the trace representation [lingsole] of QC codes) where we slightly modified the notation in the statement to better align with the notation in this paper. Moreover, if $\mathbb{F}_{q}$ is a field and $\mathbb{F}_{q^{s}}$ a a field extension of $\mathbb{F}_{q}$, the trace map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}\to\mathbb{F}_{q}$ is defined as
(12) $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a):=\sum_{i=0}^{s-1}a^{q^{i}}.$
@@
[lingsole, Theorem 5.1] Let $\mathbb{F}_{q}$ be a finite field and $\gcd(m,q)=1$. Then, for any $\ell$, the quasi-cyclic codes over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ are precisely given by the following construction. Decompose $x^{m}-1$ as per (7) . Write $\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}=F_{i}$, $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}=G_{j}^{\prime}$, and $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}=G_{j}^{\prime\prime}$. Let $U_{i}$ (resp. $V_{j}$ and $W_{j}$) denote the cyclotomic coset of $\mathbb{Z}/m\mathbb{Z}$ corresponding to $F_{i}$ (resp. $G_{j}^{\prime}$ and $G_{j}^{\prime\prime}$) and fix $u_{i}\in U_{i}$, $v_{j}\in V_{j}$, and $w_{j}\in W_{j}$. For each $i$, let $C_{i}$ be a code of length $\ell$ over $F_{i}$, and for each $j$, let $C_{j}^{\prime}$ be a code of length $\ell$ over $G_{j}^{\prime}$ and let $C_{j}^{\prime\prime}$ be a code of length $\ell$ over $G_{j}^{\prime\prime}$. For $\mathbf{x_{i}}\in C_{i}$, $\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime}$, and $\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}$, and for each $0\leq g\leq m-1$, let
(13) $c_{g}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))=\sum_{i=1}^{s}\operatorname{Tr}_{F_{i}/\mathbb{F}_{q}}(\mathbf{x_{i}}\alpha^{-gu_{i}})+\sum_{j=1}^{t}\left(\operatorname{Tr}_{G_{j}^{\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime}}\alpha^{-gv_{j}})+\operatorname{Tr}_{G_{j}^{\prime\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime\prime}}\alpha^{-gw_{j}})\right),$
(14) $\mathcal{C}=\{(c_{0}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}})),...,c_{m-1}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))):\forall\;\mathbf{x_{i}}\in C_{i},\forall\;\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime},\mbox{ and }\forall\;\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}\}$
is a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$. Conversely, every quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ is obtained through this construction.
2.2. The Duals of Quasi-Cyclic Codes
Recently, QC codes have been used in constructions of QECCs where orthogonality conditions on QC codes play a central role [Abdukhalikov-Bag-Panario2023, GalindoHernando, benjwal] . To study the duality of codes in $\mathbb{F}_{q}^{m\ell}$, we define the following inner products: Let $\textbf{a},\textbf{b}\in\mathbb{F}_{q}^{m\ell}$ where
a $=(a_{00},a_{01},...,a_{0,\ell-1},a_{10},...,a_{1,\ell-1},...,a_{m-1,0},...,a_{m-1,\ell-1})$
b $=(b_{00},b_{01},...,b_{0,\ell-1},b_{10},...,b_{1,\ell-1},...,b_{m-1,0},...,b_{m-1,\ell-1}).$
The Euclidean inner product on $\mathbb{F}_{q}^{m\ell}$ is
(15) $\mathbf{a}\cdot\mathbf{b}=\langle\textbf{a},\textbf{b}\rangle_{E}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}.,$
and the Hermitian inner product on $\mathbb{F}_{q^{t}}^{m\ell}$, for even $t$ is
(16) $\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{\sqrt{q^{t}}}.$
In particular, when $\textbf{a},\textbf{b}\in\mathbb{F}_{q^{2}}^{m\ell}$, then $\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{q}$.
On $R^{\ell}$, the Hermitian inner product is as follows: for $\textbf{x}=(x_{0},x_{1},...,x_{\ell-1})$, and $\textbf{y}=(y_{0},y_{1},...,y_{\ell-1})$,
(17) $\langle\textbf{x},\textbf{y}\rangle_{H}=\sum_{j=0}^{\ell-1}x_{j}\overline{y_{j}},$
where $\bar{\;}$ denotes the “conjugation map” on $R$ that sends $x$ to $x^{-1}=x^{m-1}$ and is extended $\mathbb{F}_{q}$-linearly. The following proposition is useful for determining when orthogonality in $R^{\ell}$ implies orthogonality in $\mathbb{F}_{q}^{m\ell}$.
Proposition 2.3.
[lingsole, Proposition 3.2] Let $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m\ell}$. Then $(T^{k\ell}(\mathbf{a}))\cdot\mathbf{b}=0$ for all $0\leq k\leq m-1$ if and only if $\langle\phi(\mathbf{a}),\phi(\mathbf{b})\rangle_{H}=0$, where $\phi$ is the aforementioned $R$-module isomorphism.
Let $\mathcal{C}$ be a QC code in $\mathbb{F}_{q}^{m\ell}$ and define the Euclidean dual, $\mathcal{C}^{\perp_{E}}$ of the code $\mathcal{C}$ by
(18) $\mathcal{C}^{\perp_{E}}:=\{\textbf{x}\in\mathbb{F}_{q}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{E}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
Similarly, the Hermitial dual, $\mathcal{C}^{\perp_{H}}$ of the code $\mathcal{C}$ is given by
(19) $\mathcal{C}^{\perp_{H}}:=\{\textbf{x}\in\mathbb{F}_{q^{t}}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{H}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
We say that $\mathcal{C}$ is self-orthogonal (ESO) with respect to the Euclidean inner product if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{E}}$, dual-containing (EDC) with respect to the Euclidean inner product if $\mathcal{C}^{\perp_{E}}\subseteq\mathcal{C}$, and self-dual (ESD) with respect to the Euclidean inner product if $\mathcal{C}^{\perp_{E}}=\mathcal{C}$. Similarly, $\mathcal{C}$ is self-orthogonal (HSO), dual-containing (HDC), and self-dual (HSD) with respect to the Hermitian inner product if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{H}},\mathcal{C}^{\perp_{H}}\subseteq\mathcal{C}$, and $\mathcal{C}=\mathcal{C}^{\perp_{H}}$, respectively.
Corollary 2.4.
[lingsole, Corollary 3.3] Let $\mathcal{C}$ be a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ and let $\phi(\mathcal{C})$ be its image in $R_{q,m}^{\ell}$ under $\phi$. Then $\phi(\mathcal{C})^{\perp}=\phi(\mathcal{C}^{\perp})$, where the dual in $\mathbb{F}_{q}^{m\ell}$ is taken with respect to the Euclidean inner product, while the dual in $R_{q,m}^{\ell}$ is taken with respect to the Hermitian inner product. In particular, a quasi-cyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if only if $\phi(\mathcal{C})$ is self-dual over $R_{q,m}$ with respect to the Hermitian inner product.
The following result, which we slightly modified to better align with the notation in this paper, allows us to characterize the duality of QC codes using constituent codes and will be used throughout this paper.
Theorem 2.5.
@@ 2.2. The Duals of Quasi-Cyclic Codes
[lingsole, Theorem 4.2] An $\ell$ quasi-cyclic code $\mathcal{C}$ of length $m\ell$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if and only if
@@
(20) $\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus(C^{\prime}_{j})^{\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}\right)$
@@
where for $1\leq i\leq s$, $C_{i}$ is a self-dual code over $F_{i}$ of length $\ell$ (with respect to the Hermitian inner product) and, for $1\leq j\leq t$, $C^{\prime}_{j}$ is a code of length $\ell$ over $G^{\prime}_{j}$ and $C_{j}^{{}^{\prime\prime}}=(C^{\prime}_{j})^{\perp_{E}}$ is its dual with respect to the Euclidean inner product.
Remark 1.
@@
    As observed in [HKS, page 136], the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all $i$. However, there are two exceptions. The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$ and we equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
@@
The next result extends [lingsole, Theorem 4.2] to the dual of a QC code.

### v4

@@ 2.2 The Cyclic Code Associated to a QC Code and a Lower Bound for its Minimum Distance
Let $s\in\mathbb{Z}_{m}$ and let $C_{s}=\{s,sq,sq^{2},...,sq^{r-1}\}\subset\mathbb{Z}_{m}$ be the $q$-cyclotomic coset of $s$, where $r$ is the smallest positive integer so that $sq^{r}\equiv s\mod m$. It is important to note that the $q$-cyclotomic cosets split $\mathbb{Z}_{m}.$ This classic concept is seen in constructions of cyclic codes, as well as in the trace representation of QC codes in  lingsole.
f $\mathbb{F}_{q}$ is a field and $\mathbb{F}_{q^{s}}$ a field extension of $\mathbb{F}_{q}$, the trace map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}\to\mathbb{F}_{q}$ is defined as
$\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a):=\sum_{i=0}^{s-1}a^{q^{i}},$
and we can extend the trace map componentwise to $\mathbb{F}_{q^{s}}^{\ell}$ by defining the $\mathbb{F}_{q}$-linear map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}^{\ell}\rightarrow\mathbb{F}_{q}^{\ell}$ as
$\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a_{1},\dots,a_{\ell})=\big(\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a_{1}),\dots,\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}(a_{\ell})\big).$
theorem(lingsole, Theorem 5.1) Let $\mathbb{F}_{q}$ be a finite field and $\gcd(m,q)=1$. Then, for any $\ell$, the quasi-cyclic codes over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ are precisely given by the following construction. Decompose $x^{m}-1$ as per (5) . Write $\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}=F_{i}$, $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}=G_{j}^{\prime}$, and $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}=G_{j}^{\prime\prime}$. Let $U_{i}$ (resp. $V_{j}$ and $W_{j}$) denote the cyclotomic coset of $\mathbb{Z}/m\mathbb{Z}$ corresponding to $F_{i}$ (resp. $G_{j}^{\prime}$ and $G_{j}^{\prime\prime}$) and fix $u_{i}\in U_{i}$, $v_{j}\in V_{j}$, and $w_{j}\in W_{j}$. For each $i$, let $C_{i}$ be a code of length $\ell$ over $F_{i}$, and for each $j$, let $C_{j}^{\prime}$ be a code of length $\ell$ over $G_{j}^{\prime}$ and let $C_{j}^{\prime\prime}$ be a code of length $\ell$ over $G_{j}^{\prime\prime}$. For $\mathbf{x_{i}}\in C_{i}$, $\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime}$, and $\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}$, and for each $0\leq g\leq m-1$, let
$c_{g}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))=\sum_{i=1}^{s}\operatorname{Tr}_{F_{i}/\mathbb{F}_{q}}(\mathbf{x_{i}}\alpha^{-gu_{i}})+\sum_{j=1}^{t}\left(\operatorname{Tr}_{G_{j}^{\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime}}\alpha^{-gv_{j}})+\operatorname{Tr}_{G_{j}^{\prime\prime}/\mathbb{F}_{q}}(\mathbf{y_{j}^{\prime\prime}}\alpha^{-gw_{j}})\right),$ (11)
$\mathcal{C}=\{(c_{0}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}})),...,c_{m-1}((\mathbf{x_{i}}),(\mathbf{y_{j}^{\prime}}),(\mathbf{y_{j}^{\prime\prime}}))):\forall\;\mathbf{x_{i}}\in C_{i},\forall\;\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime},\mbox{ and }\forall\;\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}\}$ (12)
is a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$. Conversely, every quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ is obtained through this construction. Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of an $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By (cem, Lemma 4.1), the trace-formula described in (lingsole, Theorem 5.1) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, when taking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of an $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
$BZ(D^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i=1,2,...,h\}.$
Recall that if $D$ is a cyclic code of length $m$ over $\mathbb{F}_{q}$ and $\alpha$ is a primitive $m$-th root of unity, the basic zero set of $D^{\perp}$ consists of those powers $\alpha^{-u}$ at which every codeword polynomial of $D^{\perp}$ vanishes. Equivalently, it is the set of roots of the generator polynomial of $D^{\perp}$, described via representatives of their $q$-cyclotomic cosets modulo $m$. Just as in BastosAlvarezFloresSalerno, we call $D$ the cyclic code associated to $\mathcal{C}$. Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
$1\leq i_{1}<i_{2}<\cdots<i_{t}\leq h.$
Define $D_{I}:=D_{i_{1},...,i_{t}}\subset D$ to be the cyclic subcode of $D$ whose dual’s basic zero set is
$BZ(D_{I}^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i\in I\}$
so that $D_{I}=D_{i_{1}}\oplus D_{i_{2}}\oplus\cdots\oplus D_{i_{t}}$ and $d(D_{I})\geq d(D)$ for each $I\subset\{1,2,...,h\}$.
gray[AA: I moved this gray text up to the section 2.3] Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of an $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By (cem, Lemma 4.1), the trace-formula described in (11) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, looking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of an $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
$BZ(D^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i=1,2,...,h\}.$
[As per reviewer, include an explanation on the “basic zero set"] Recall that if $D$ is a cyclic code of length $m$ over $\mathbb{F}_{q}$ and $\alpha$ is a primitive $m$-th root of unity, the basic zero set of $D^{\perp}$ consists of those powers $\alpha^{-u}$ at which every codeword polynomial of $D^{\perp}$ vanishes. Equivalently, it is the set of roots of the generator polynomial of $D^{\perp}$, described via representatives of their $q$-cyclotomic cosets modulo $m$. Just as in  BastosAlvarezFloresSalerno, the we call $D$ the cyclic code associated to $\mathcal{C}$. In this section, we set $D_{i}=\langle\theta_{i}\rangle$, where $BZ(D_{i}^{\perp})=\{\alpha^{-u_{i}}\}$ for each $1\leq i\leq h$. Note that $D=D_{1}\oplus D_{2},\oplus\cdots D_{h}$, and that each $D_{i}$ is a subcode of $D$. Hence, $d(D_{i})\geq d(D)$ for each $1\leq i\leq h$.
Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
$1\leq i_{1}<i_{2}<\cdots<i_{t}\leq h.$
Define $D_{I}:=D_{i_{1},...,i_{t}}\subset D$ to be the cyclic subcode of $D$ whose dual’s basic zero set is
$BZ(D_{I}^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i\in I\}$ (13)
such that $D_{I}=D_{i_{1}}\oplus D_{i_{2}}\oplus\cdots\oplus D_{i_{t}}$.
efine the following quantity:
$R_{I}(\mathcal{C})$ $=R_{i_{1},...,i_{t}}(\mathcal{C})$ (14)
$:=\begin{cases}d(C_{i_{1}})d(D_{i_{1}})&\text{if $t=1$}\\ \\ \left(d(C_{i_{1}})-d(C_{i_{2}})\right)d(D_{i_{1}})+\left(d(C_{i_{2}})-d(C_{i_{3}})\right)d(D_{i_{1},i_{2}})+\\ \hskip 128.0374pt\vdots&\text{if $t\geq 2$}\\ +\left(d(C_{i_{t-1}})-d(C_{i_{t}})\right)d(D_{i_{1},i_{2},...i_{t-1}})+d(C_{i_{t}})d(D_{i_{1},i_{2},...,i_{t}}).\end{cases}$ (15)
The following result uses Jensen’s concatenated description of QC codes (OO24, Theorem 2.1), and their associated cyclic codes and subcodes to establish a lower bound on the minimum distance of a QC code. This lower bound will be referred to Jensen’s bound where relevant details can be be found in jensen; OO24.
Theorem 4
(jensen, Theorem 4) Given a QC code $\mathcal{C}$ of length $m\ell$ and index $\ell$ over $\mathbb{F}_{q}$ with the concatenated structure ${\mathcal{C}\cong\bigoplus_{i=1}^{h}D_{i}\square C_{i}},$ and $d(C_{1})\geq d(C_{2})\geq...\geq d(C_{h})$, we have
$d(\mathcal{C})\geq d_{\text{J}}(\mathcal{C}):=\min\{N_{h},N_{h-1,h},\ldots,N_{1,2,\ldots,h}\},$ (16)
where
$N_{h}(\mathcal{C})$ $=d(C_{h})d(D_{h}),$
$N_{h-1,h}(\mathcal{C})$ $=d(C_{h-1})d(D_{h-1,h}),$
$\vdots$
$N_{1,2,\ldots,h}(\mathcal{C})$ $=d(C_{1})d(D_{1,2,\ldots,h}).$
Using the CRT decomposition (9) and Jensen’s concatenated description, we know that a QC code $\mathcal{C}$ can be decomposed as the following direct sum of concatenations
$\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}\left(D_{j}^{\prime}\square C_{j}^{\prime}\oplus D_{j}^{\prime\prime}\square C_{j}^{\prime\prime}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}D_{i}\square C_{i}\right).$ (17)
We will apply Theorem 4 to the QC code representation in (17) to determine a lower bound for the minimum distance for the codes in our QC family in Section 3.
theorem[Güneri-Özbudak bound](cem, Theorem 4.8) Let $\mathcal{C}$ be a QC code as in (9) with the assumption (10). Then we have
$d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{R_{h}(\mathcal{C}),R_{h-1,h}(\mathcal{C}),R_{h-2,h-1,h}(\mathcal{C}),...,R_{1,2,...,h}(\mathcal{C})\}.$ (18)
lightgray
Remark 2.5.
  1. (i)
    If the number of nonzero constituent codes is two, the Jensen bound (4) becomes
    $d(\mathcal{C})\geq d_{\text{J}}(\mathcal{C}):=\min\{d(C_{2})d(D_{2}),d(C_{1})d(D)\},$
  2. (ii)
    If the number of nonzero constituent codes is three, the Jensen bound (4) becomes
    $d(\mathcal{C})\geq d_{\text{J}}(\mathcal{C}):=\min\{N_{3},N_{2,3},N_{1,2,3}\},$
    where $N_{3}=d(C_{3})d(D_{3}),N_{2,3}=d(C_{2})d(D_{2,3})$ and $N_{1,2,3}=d(C_{1})d(D)$.
2.3 The Duals of Quasi-Cyclic Codes
Recently, QC codes have been used in constructions of QECCs where orthogonality conditions on QC codes play a central role Abdukhalikov-Bag-Panario2023; GalindoHernando; benjwal . To study the duality of codes in $\mathbb{F}_{q}^{m\ell}$, we define the following inner products: Let $\textbf{a},\textbf{b}\in\mathbb{F}_{q}^{m\ell}$ where $\textbf{a}=(a_{ij})$ and $\textbf{b}=(b_{ij})$ where $0\leq i\leq m-1$ and $0\leq j\leq\ell-1$. align* a =(a_00, a_01,…,a_0,ℓ-1, a_10,…, a_1, ℓ-1,…,a_m-1,0,…,a_m-1, ℓ-1)
b =(b_00, b_01,…,b_0,ℓ-1, b_10,…, b_1, ℓ-1,…,b_m-1,0,…,b_m-1, ℓ-1). The Euclidean inner product on $\mathbb{F}_{q}^{m\ell}$ is
$\mathbf{a}\cdot\mathbf{b}=\langle\textbf{a},\textbf{b}\rangle_{E}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij},$
and the Hermitian inner product on $\mathbb{F}_{q^{t}}^{m\ell}$, for even $t$ is
$\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{\sqrt{q^{t}}}.$
In particular, when $\textbf{a},\textbf{b}\in\mathbb{F}_{q^{2}}^{m\ell}$, then $\langle\textbf{a},\textbf{b}\rangle_{H}=\sum_{i=0}^{m-1}\sum_{j=0}^{\ell-1}a_{ij}b_{ij}^{q}$. On $R_{q,m}^{\ell}$, the Hermitian inner product is as follows: for $\textbf{x}=(x_{0},x_{1},...,x_{\ell-1})$, and $\textbf{y}=(y_{0},y_{1},...,y_{\ell-1})$,
$\langle\textbf{x},\textbf{y}\rangle_{H}=\sum_{j=0}^{\ell-1}x_{j}\overline{y_{j}},$
where $\bar{\;}$ denotes the “conjugation map" on $R_{q,m}$ that sends $x$ to $x^{-1}=x^{m-1}$ and is extended $\mathbb{F}_{q}$-linearly. The following proposition is useful for determining when orthogonality in $R_{q,m}^{\ell}$ implies orthogonality in $\mathbb{F}_{q}^{m\ell}$.
Proposition 2.6.
(lingsole, Proposition 3.2) Let $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m\ell}$. Then $(T^{k\ell}(\mathbf{a}))\cdot\mathbf{b}=0$ for all $0\leq k\leq m-1$ if and only if $\langle\phi(\mathbf{a}),\phi(\mathbf{b})\rangle_{H}=0$, where $\phi$ is the aforementioned $R_{q,m}$-module isomorphism.
Let $\mathcal{C}$ be a QC code in $\mathbb{F}_{q}^{m\ell}$ and define its Euclidean dual $\mathcal{C}^{\perp_{E}}$ by
$\mathcal{C}^{\perp_{E}}:=\{\textbf{x}\in\mathbb{F}_{q}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{E}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
Similarly, the Hermitial dual, $\mathcal{C}^{\perp_{H}}$ of $\mathcal{C}$ is given by
$\mathcal{C}^{\perp_{H}}:=\{\textbf{x}\in\mathbb{F}_{q^{t}}^{m\ell}\mid\langle\textbf{x},\textbf{c}\rangle_{H}=0,\forall\;\textbf{c}\in\mathcal{C}\}.$
We say that $\mathcal{C}$ is self-orthogonal with respect to the Euclidean inner product (or ESO) if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{E}}$, dual-containing with respect to the Euclidean inner product (or EDC) if $\mathcal{C}^{\perp_{E}}\subseteq\mathcal{C}$, and self-dual with respect to the Euclidean inner product (or ESD) if $\mathcal{C}^{\perp_{E}}=\mathcal{C}$. Analogously, $\mathcal{C}$ is self-orthogonal (HSO), dual-containing (HDC), and self-dual (HSD) with respect to the Hermitian inner product if $\mathcal{C}\subseteq\mathcal{C}^{\perp_{H}},\mathcal{C}^{\perp_{H}}\subseteq\mathcal{C}$, and $\mathcal{C}=\mathcal{C}^{\perp_{H}}$, respectively.
lightgray
Corollary 2.7.
(lingsole, Corollary 3.3) Let $\mathcal{C}$ be a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ and let $\phi(\mathcal{C})$ be its image in $R_{q,m}^{\ell}$ under $\phi$. Then $\phi(\mathcal{C})^{\perp}=\phi(\mathcal{C}^{\perp})$, where the dual in $\mathbb{F}_{q}^{m\ell}$ is taken with respect to the Euclidean inner product, while the dual in $R_{q,m}^{\ell}$ is taken with respect to the Hermitian inner product. In particular, a quasi-cyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if only if $\phi(\mathcal{C})$ is self-dual over $R_{q,m}$ with respect to the Hermitian inner product.
The following result, which we slightly modified to better align with the notation in this paper, allows us to characterize the duality of QC codes using constituent codes and will be used throughout this paper to determine the duality of the QC codes in our code construction in Section 3.
@@ 2.3 The Duals of Quasi-Cyclic Codes
(lingsole, Theorem 4.2) An $\ell$-quasi-cyclic code $\mathcal{C}$ of length $m\ell$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if and only if
@@
$\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus(C^{\prime}_{j})^{\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}\right)$ (19)
@@
where for $1\leq i\leq s$, $C_{i}$ is a self-dual code over $F_{i}$ of length $\ell$ (with respect to the Hermitian inner product) and, for $1\leq j\leq t$, $C^{\prime}_{j}$ is a code of length $\ell$ over $G^{\prime}_{j}$ and $C_{j}^{{}^{\prime\prime}}=(C^{\prime}_{j})^{\perp_{E}}$ is its dual with respect to the Euclidean inner product over $G_{j}^{{}^{\prime\prime}}$.
Remark 2.9.
@@
    As observed in (HKS, page 136), the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all $i$. However, there are two exceptions: The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$, so we can equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
@@
lightgray Thus, when we apply Theorem 2.8 to Jensen’s concatenated description in (17), we obtain
$\mathcal{C}\cong\left(\bigoplus_{j=1}^{t}\left((D_{j}^{\prime}\square C_{j}^{\prime})\oplus(D_{j}^{\prime}\square C_{j}^{\prime})^{\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}D_{i}\square C_{i}\right).$ (20)
Such representation of a QC code $\mathcal{C}$ will be vital when computing a lower bound for the minimum distance of $\mathcal{C}$ in Section 3. The next result extends (lingsole, Theorem 4.2) to the dual of a QC code.